# Closing the Data-Efficiency Gap Between Autoregressive and Masked Diffusion LLMs

### v2

@@ Abstract
Despite autoregressive large language models (arLLMs) being the current dominant paradigm in language modeling, effectively updating these models to incorporate new factual knowledge still remains difficult. They resist knowledge injection via fine-tuning due to inherent shortcomings such as the “reversal curse” — the challenge of answering questions that reverse the original information order in the training sample. Masked diffusion large language models (dLLMs) are rapidly emerging as a powerful alternative to the arLLM paradigm, with evidence of better data efficiency and free of the “reversal curse” in pre-training. However, it is unknown whether these advantages extend to the post-training phase, i.e. whether pre-trained dLLMs can easily acquire new knowledge through fine-tuning. On three diverse datasets, we fine-tune arLLMs and dLLMs, evaluating them with forward and backward style Question Answering (QA) to probe knowledge generalization and the reversal curse. Our results confirm that arLLMs critically rely on extensive data augmentation via paraphrases for QA generalization, and paraphrases are only effective when their information order matches the QA style. Conversely, dLLMs achieve high accuracies on both forward and backward QAs without paraphrases; adding paraphrases yields only marginal gains. Inspired by the dLLM’s performance, we introduce a novel masked fine-tuning paradigm for knowledge injection into pre-trained arLLMs. This proposed method successfully and drastically improves the data efficiency of arLLM fine-tuning, effectively closing its performance gap with dLLMs. We further show that the masked fine-tuning paradigm of arLLMs can be extended to the supervised fine-tuning (SFT) of mathematical capability. Across two models and two datasets, our masked SFT outperforms regular SFT.
Figure 1: A schematic summary of the results. First row: autoregressive LLM requires paraphrases for generalizing knowledge in the fine-tuning text to QA tasks, and suffer from reversal curse (i.e. fail to answer backward questions). Second row: masked diffusion LLM can easily generalize fine-tuning text to QA tasks in both forward and backward styles. Third row: inspired by the masked diffusion LLM, we propose a masked fine-tuning paradigm, that closes the fine-tuning gap between autoregressive LLMs and masked diffusion LLMs.
@@ 1 Introduction
Despite auto-regressive large language models (arLLMs) being the main contributors to the modern success of language modeling, studies have demonstrated the difficulty of injecting new knowledge into pre-trained arLLMs by fine-tuning on documents that are not in the pre-training dataset (ovadia2023fine; mecklenburg2024injecting; gekhman2024does; soudani2024fine; zhao2025style; lampinen2025generalization). Fine-tuned models typically generalize poorly to downstream tasks such as question-answering (QA). An example failure mode is the famous “reversal curse”, that LLMs fail to answer the questions in the reversed order of the training text (berglund2023reversal). Fine-tuning on multiple rewrites (i.e., paraphrases) of the documents can mitigate generalization issues but still falls behind in-context learning based external memory systems like RAG (ovadia2023fine; mecklenburg2024injecting). This pitfall of arLLMs is a major obstacle that limits current models from being flexible life-long learners via weight updates.
As alternatives to the auto-regressive models, several recent masked diffusion large language models (dLLMs) have been scaled up to be as capable as arLLMs on multiple downstream tasks, with additional advantages such as high-throughput decoding of multiple tokens simultaneously (ICLR2025_ce1c1ff5; llada; ye2025dream). Instead of factorizing the joint sequence probability strictly left-to-right, dLLMs learn the factorization under many permutations of token positions, enabling them to condition on any subset of tokens and predict the remainder. Though such an objective is harder than learning autoregressive factorization (kim2025train), dLLMs do not suffer from “reversal curse(llada), and can achieve lower validation loss (i.e. token cross-entropy loss) than arLLMs in data-constrained settings (prabhudesai2025diffusion; ni2025difflm). However, most of the dLLMs studies have focused on the properties in the pre-training phase. Little is known about whether dLLMs also have advantages in the post-training phase, such as knowledge injection by fine-tuning.
In this study, we use three datasets to compare the data efficiency and performance of knowledge injection by fine-tuning in arLLM and dLLM models. We refer to “data-efficiency” as how much fine-tuning data (including paraphrases) is required for a pre-trained model to generalize the knowledge in the data to downstream tasks such as QA. We introduce a novel masked fine-tuning paradigm for arLLMs that emulates a diffusion-style mask reconstruction loss. Across all datasets, dLLMs show a consistent data-efficiency advantage over arLLMs, and our masked fine-tuning largely closes this gap, bringing arLLMs to strong performance without relying on paraphrases. We further show that adapting demasking objection to supervised fine-tuning (SFT) of arLLM outperforms regular SFT on math tasks. More specifically, we show the following results:
  • arLLMs heavily rely on paraphrases to successfully generalize fine-tuning text to downstream QA tasks; arLLMs fail on backward style questions, and only paraphrases that reverse the information order in the sentences can mitigate the reversal curse.
  • dLLMs can achieve high accuracy in both forward and backward questions without paraphrases; adding paraphrases only marginally helps. This establishes the knowledge injection data efficiency of dLLMs in the post-training phase.
  • We propose a masked fine-tuning paradigm that fine-tunes arLLMs in a “masked infilling” way by giving masked samples in the context with instructions to recover the mask, and set the unmasked sample as the supervised fine-tuning target. The novel method closes the performance gap between arLLMs and dLLMs fine-tuning: it allows arLLMs of different sizes and model families to achieve strong performance in both forward and backward questions without paraphrases.
  • We extend the masked fine-tuning paradigm to SFT tasks, and show that across two arLLMs and two math datasets masked SFT outperforms regular SFT.
Taken together, these results show that dLLMs are more data-efficient than arLLMs during post-training. We further show that this advantage can be transferred to arLLMs via our masked fine-tuning paradigm. Our findings suggest the possibility to post-train an LLM to adapt to the changing world using a small amount of new knowledge texts, which could help address challenges in keeping AI systems updated with changing environment.
@@ 2 Background
2.1 Knowledge injection by fine-tuning and Reversal curse
An AI system intended for real-world deployment should continually acquire and integrate new information to adapt to evolving environments. Though LLMs have been successful on numerous tasks, they struggle to incorporate new knowledge into their weights. At least two factors contribute to this challenge. First, fine-tuning on new tasks can induce catastrophic forgetting of previously learned capabilities (luo2023empirical; wang2023trace; zhai2023investigating; zhang2024dissecting; chen2024chatgpt; ren2024analyzing). Second, fine-tuning primarily reshapes surface behavior (e.g., tone and format) without reliably integrating new factual knowledge (ovadia2023fine; mecklenburg2024injecting; gekhman2024does; soudani2024fine; zhao2025style; lampinen2025generalization).
@@ 2.1 Knowledge injection by fine-tuning and Reversal curse
A failure mode of learning knowledge in the text is the “reversal curse”, that after learning statements of the form “A is B”, the model does not generalize to its inverse form “B is A”. The reversal curse has been observed across the training phases and models (berglund2023reversal; zhu24b; lv-etal-2024-analysis; lin2024delving; guo-etal-2024-mitigating; golovneva2024reverse; lu2024rethinking). Even strong commercial models like GPT-4 and GPT-4o show signs of the reversal curse (berglund2023reversal; llada). The cause of the reversal curse has been theoretically attributed to an inherent limitation of the autoregressive training objective (zhu2024towards; kitouni2024factorization) (more discussion in Appendix A.6). Common approaches to mitigate the reversal curse in autoregressive models include: (i) augmenting the training set with paraphrases (lu2024rethinking), which requires substantial computation to construct; (ii) augmenting the training set with reordered sequences (guo-etal-2024-mitigating; golovneva2024reverse), which often violate natural-language grammar and degrade overall language modeling performance; and (iii) replacing causal attention with bidirectional attention (lv-etal-2024-analysis), which struggles to retrieve long information (e.g., a person’s description). In contrast, our masked fine-tuning paradigm for arLLMs addresses the reversal curse without constructing paraphrase augmentations or altering the autoregressive objective. With only adjustments to the fine-tuning query and keeping the original documents as optimization targets, our method attains high accuracy across all datasets we evaluated.
2.2 Masked diffusion language models
@@
Recently, dLLMs have emerged as a strong competitor to arLLMs (sahoo2024simple; llada; ye2025dream). Comparing to autoregressive models, dLLMs use encoder-only transformers to generate text by iteratively unmasking tokens via a reversed discrete diffusion process. The training objective is to minimize the mask reconstruction loss llada:
@@
$\mathcal{L}(\theta)=-\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{t}}\left[\frac{1}{t}\sum_{\ell=1}^{L}\mathbb{I}[\bm{x}_{t}^{\ell}\in\text{M}]\log p_{\theta}(\bm{x}_{0}^{\ell}|\bm{x}_{t})\right],$ (1)
@@
$\bm{x_{0}}$ is an original sequence of length $L$ sampled from the training data. The masking process is governed by the mask ratio $t$, which is sampled uniformly, resulting in the corrupted sequence $\bm{x_{t}}$. The set $\mathbf{M}$ denotes the indices of the tokens that were masked by the forward process at ratio $t$. The $\ell$-th token is considered for the loss only if it was masked. Such a loss objective has been shown to be the negative evidence lower bound (ELBO) on the data likelihood (shi2024simplified).
@@
Recent studies report dLLMs are more data-efficient than arLLMs. When the training data is scarce, dLLMs keep improving with repeated use of the data and surpass arLLMs on validation loss, while arLLMs saturate the validation loss or increase it due to overfitting (prabhudesai2025diffusion; ni2025difflm). prabhudesai2025diffusion further shows that the lower validation loss in dLLMs can generalize to downstream tasks like ARC-Easy, and attributes its data efficiency to random masks as implicit data augmentation. However, whether these advantages persist in new knowledge acquisition during post-training—where the model needs to learn generalizable knowledge through small fine-tuning sets—remains unclear.
Note that the denoising objective and bi-directional attention of dLLMs is closely related to the design choices in earlier language models such as BERT and T5/Flan-T5 devlin2019bert; raffel2020exploring. However, they are less popular choices for studying the post-training phase properties. BERT is not a generative chat models thus cannot achieve QA tasks; the complexity of the encoder-decoder design of T5 hinders its advancement to match other state-of-art LLMs. Thus, in this study, we choose to use state-of-the-art dLLMs and arLLMs to probe the post-training properties involving question-answering.
3 Datasets and experimental setups
We focus on assessing LLMs’ ability to learn new knowledge through fine-tuning. More specifically, LLMs are fine-tuned on a set of documents that contain knowledge unknown to the base LLM, and evaluated by open-ended QA tasks. The correctness of an answer is evaluated by the ROUGE-1 score (lin2024delving; jiang2025anyedit) between the generated answer and the ground truth answer, which we report as “accuracy.” It measures the proportion of the words in the ground truth answer that appear in the generated answer. To better demonstrate the generation quality, we also show examples of model responses in all the experiments in A.7.
In the first part of the study, we use three representative datasets. Two are existing synthetic datasets from previous studies on the reversal curse, and each accompanied with a paraphrase set as data augmentation; one is constructed using recent Wikipedia articles. Previous studies on the reversal curse rarely use realistic datasets. In this study, we include the Wiki dataset to probe the reversal curse in a setting that reflects knowledge injection in the real world. See examples of each dataset in Appendix A.3.
The NameDescription (berglund2023reversal) contains 60 statements of different fictitious individuals, 30 each of the form “[name] is [description]” (N2D) and “[description] is [name]” (D2N). lin2024delving extended the dataset with an open-ended QA testing set. For each type of statement, the QA set contains two types of questions: “What is the name related to a given description” and “What is the description of a given name”. Depending on whether the question is aligned with the original statement, each question is classified as “forward” or “backward” question (e.g. N2D statement with “What is the description of a given name” type of question is a forward question). The dataset also contains a paraphrase set, in which each statement is rewritten into 30 different versions, but the order of [name] and [description] in the paraphrases is always preserved as in the original statement (either N2D or D2N).
The Biography dataset is proposed in zhu24a; zhu24b. Since the original dataset is not publicly available, we used a subset of 100 samples from a replication (zheng2025spurious). Each sample is a 6-sentence paragraph about a fictitious individual, detailing their birth city, birthday, college, and job information. Note that the name only appears in the first sentence and is replaced with a pronoun in the following sentences; thus, questions about the name are considered backward questions. Each sample also includes a paraphrase set of 5 paraphrases; the paraphrases do not change the order of the sentences but only alter the wording while preserving the information. The testing QA set has both forward (i.e., asking for an attribute given the name) and backward styles (i.e., asking for the name given 3 attributes from the person) questions.
The Wiki dataset contains 94 Wikipedia articles, constructed according to the protocol described in pan2025memorization. We crawl the Wikipedia pages under the category “2025 by month”, and then further filter out the pages that were created before the year 2025. This procedure ensures that these real-world events are recent enough that both dLLM and arLLM models have minimal knowledge about them, which is justified by the models’ accuracy before fine-tuning (Table 1). For each wiki article, we use GPT-o3-mini to generate QA pairs in both forward and backward styles. By prompting GPT-o3-mini, we construct two different paraphrase sets: one that retains the information in place while only changing the wording (same-order paraphrases); the other also changes the order of information in the article (permute-order paraphrases). 10 paraphrases of each type are generated for every wiki article. All LLM generated QAs or paraphrases are cross-checked by human and invalid entries are filtered out and replaced. More details on constructing the datasets are provided in Appendix A.3.
We chose the arLLMs Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (dubey2024llama), Qwen2.5-7B-Instruct, Qwen3-4B-Instruct-2507 yang2025qwen3 and dLLM LLaDA-8B-Instruct (llada) to conduct the experiments. LLaDA-8B-Instruct is directly comparable to Llama-3.1-8B-Instruct as they have similar parameter size and comparable general capability benchmarks. Fine-tuning and evaluation configurations are provided in the Appendix A.4.
@@ 3 Datasets and experimental setups
Masked Llama 8B w/o paraphrases 0.658 0.949 0.992 0.923 0.971 0.598 0.980 0.930
Masked Llama 8B w paraphrases 0.969 0.996 0.928 0.832 0.965 0.816 0.905 0.794
@@
Masked Llama 3B w/o paraphrases 0.887 0.932 0.992 0.933 0.967 0.738 0.970 0.908
Masked Llama 3B w paraphrases 0.982 0.928 1.000 0.992 0.964 0.809 0.855 0.810
@@
Masked Qwen 7B w/o paraphrases 0.870 0.897 0.967 0.929 0.984 0.754 0.934 0.896
Masked Qwen 7B w paraphrases 0.957 0.810 0.933 1.000 0.960 0.828 0.867 0.821
@@
Masked Qwen 4B w/o paraphrases 0.928 0.902 0.950 0.967 0.944 0.623 0.907 0.870
Masked Qwen 4B w paraphrases 0.975 0.816 1.000 0.975 0.967 0.806 0.847 0.822
Table 1: Fine-tuning performance of the dLLM and arLLMs across all three datasets. Model names are shortened for presentation (see section 3 for full model names). “Masked” denotes the masked fine-tuning paradigm. “Reverse training” denotes the entity-level reverse-training paradigm proposed in golovneva2024reverse, which we use as a baseline control. For the Wiki dataset, we use the same-order paraphrase set. Pink indicates clear failure (accuracy below 50%); turquoise indicates clear success (accuracy above 90%).
4 arLLM knowledge injection relies on paraphrases

### v3

# Diffusion-Inspired Masked Fine-Tuning for Knowledge Injection in Autoregressive LLMs

@@ Abstract
Large language models (LLMs) are often used in environments where facts evolve, yet factual knowledge updates via fine-tuning on unstructured text often suffers from 1) reliance on compute-heavy paraphrase augmentation and 2) the reversal curse. Recent studies show diffusion large language models (dLLMs) require fewer training samples to achieve lower loss in pre-training and are more resistant to the reversal curse, suggesting dLLMs may learn new knowledge more easily than autoregressive LLMs (arLLMs). We test this hypothesis in controlled knowledge fine-tuning experiments and find that while arLLMs rely on paraphrase augmentation to generalize knowledge text into question-answering (QA) capability, dLLMs do not require paraphrases to achieve high QA accuracy. To further investigate whether the demasking objective alone can induce such a knowledge injection advantage in dLLMs regardless of their diffusion denoising paradigm, we propose masked fine-tuning for arLLMs, which prompts an arLLM to reconstruct the original text given a masked version in context. The masked fine-tuning for arLLMs substantially improves the efficacy of knowledge injection, i.e. no paraphrase needed and resistant to the reversal curse, closing the gap between arLLMs and dLLMs. We also demonstrate that the same demasking objective improves supervised fine-tuning (SFT) on math tasks over standard SFT, suggesting broader applicability of the demasking objective.
Machine Learning, ICML
Figure 1: A schematic summary of the results. First row: autoregressive LLM requires paraphrases to generalize knowledge from the fine-tuning text to QA tasks, and suffers from reversal curse (i.e. fail to answer backward questions). Second row: masked diffusion LLM can easily generalize fine-tuning text to QA tasks in both forward and backward styles. Third row: Inspired by the masked diffusion LLM, we propose a masked fine-tuning paradigm that closes the fine-tuning gap between autoregressive LLMs and masked diffusion LLMs.
@@ 1 Introduction
Large language models (LLMs) as general assistants are often deployed in settings where facts evolve: news breaks, policies change, and organizations maintain internal knowledge that is continuously updated. A natural idea to keep an LLM up-to-date is to fine-tune the model on newly available documents, in a way similar to pre-training. Yet, a growing body of work shows that fine-tuning on knowledge documents often struggles to translate into reliable downstream question answering (QA) ability, even when the fine-tuning loss decreases (Ovadia et al., 2023; Mecklenburg et al., 2024; Gekhman et al., 2024; Soudani et al., 2024; Zhao et al., 2025; Lampinen et al., 2025). This gap limits the practicality of weight updating as a long-term memory mechanism in LLMs.
There are two empirical obstacles that make knowledge injection via fine-tuning in autoregressive LLMs (arLLMs) difficult. (1) arLLMs often fail to generalize from raw documents to QA behavior unless the fine-tuning data is expanded with many paraphrases (Ovadia et al., 2023; Mecklenburg et al., 2024). This kind of augmentation is expensive and sometimes impractical, as high-quality paraphrases typically require additional LLM calls, careful filtering, and substantial engineering to avoid distribution shift. Such paraphrase dependence indicates that standard fine-tuning may not provide enough training signal from each example. (2) After training on statements whose information is presented in one order (e.g., “A is B”), arLLMs can answer questions aligned with that order but fail catastrophically when asked to invert it (e.g., “B is A”), a behavior known as the reversal curse (Berglund et al., 2023; Allen-Zhu and Li, 2025; Lv et al., 2024; Lin et al., 2024; Guo et al., 2024; Golovneva et al., 2024a; Lu et al., 2024). The reversal curse is not just “lack of data”, whether augmentation helps depends on whether it re-expresses the same facts in the order demanded by the question.
Recent masked diffusion language models (dLLMs) (Nie et al., 2025a, b; Ye et al., 2025) have revived a bidirectional “denoising” style objective, such as in BERT and T5/Flan-T5 (Devlin et al., 2019; Raffel et al., 2020), at LLM scale. Several advantages of dLLMs have been found over arLLMs. During pretraining, dLLMs can reach lower training loss in data-scarce regimes (Prabhudesai et al., 2025; Ni et al., 2025), suggesting better sample efficiency. After instruction tuning, dLLMs appear less prone to the “reversal curse”; an 8B dLLM has been reported to outperform GPT-4o on a reversal poem completion task (Nie et al., 2025b). This motivates our hypothesis about dLLM knowledge injection via fine-tuning:
Hypothesis 1: Unlike arLLMs, which depend on paraphrase augmentation and exhibit the reversal curse, dLLMs can generalize from fine-tuning on the knowledge text alone to strong QA performance, without additional paraphrased data.
We test this hypothesis in controlled knowledge injection settings. Across datasets, we find that arLLMs strongly depend on paraphrase augmentation for forward QA generalization and still fail on backward QA (an indication of the reversal curse), while dLLMs achieve high accuracy in both directions without paraphrases.
Since dLLMs’ training repeatedly asks the model to reconstruct missing tokens given a partially observed context, this gives the model many distinct conditioning patterns for the same underlying sequence, which can be viewed as implicit augmentation of token order and context availability. This motivates our second hypothesis:
Hypothesis 2: The knowledge-injection advantage of dLLMs is mainly due to the demasking training objective, rather than diffusion-style decoding or architectural differences. Therefore, adapting a demasking-style fine-tuning objective to decoder-only arLLMs should substantially improve knowledge injection from raw text and largely close the performance gap to dLLMs.
To test this hypothesis and attempt to transfer dLLM’s advantages to arLLM fine-tuning, we propose masked fine-tuning for arLLMs, a simple paradigm that reframes each knowledge document as a reconstruction task. During fine-tuning, we create a masked version of the document, prompt the model to recover the masked passage, and supervise the model on the original unmasked text as the target. By sampling different masks across steps, a single document yields many conditioning patterns, providing bidirectional learning signals while preserving the model’s decoder-only architecture and standard autoregressive training. Empirically, this masked fine-tuning objective substantially improves knowledge injection: it reduces reliance on paraphrase augmentation for forward QA and mitigates the reversal curse for backward QA, largely closing the gap to dLLMs. To probe whether the benefit is specific to factual knowledge injection or also good for procedural knowledge learning, we adapt the demasking formulation to math SFT by masking parts of the target solution and training the model to reconstruct them. Across two math datasets and two arLLMs, masked SFT improves over standard SFT, suggesting a broader applicability of the demasking objective.
@@ 2 Background
2.1 Knowledge injection and paraphrase dependence
Empirically, LLMs can appear to store facts yet fail to extract them under novel question wordings unless the training samples cover a variety of ways that the same fact can be presented. In a synthetic setting, Allen-Zhu and Li (2024) found that reliable knowledge extraction correlates strongly with paraphrase augmentation during training, and that post-hoc instruction tuning cannot fully recover extractability if the pretraining signal lacks such variation. In a more realistic setting involving Wikipedia knowledge injection into instruction-tuned LLMs, Ovadia et al. (2023) and Mecklenburg et al. (2024) found that QA capability scales with the number of paraphrases used in training but saturates quickly at around 10 paraphrases per sample and still lags behind retrieval-based methods such as RAG. Ovadia et al. (2025) showed that a multi-stage data augmentation pipeline—first breaking articles down into individual concepts and facts and then paraphrasing—can further improve QA accuracy, but such methods are extremely expensive and rely on a strong LLM to prepare the data.
2.2 Reversal curse
@@
The reversal curse describes a failure mode in LLM training in which, after learning statements of the form “A is B,” a model does not generalize to the inverse form “B is A.” The reversal curse has been observed across training phases and model families (Berglund et al., 2023; Allen-Zhu and Li, 2025; Lv et al., 2024; Lin et al., 2024; Guo et al., 2024; Golovneva et al., 2024a; Lu et al., 2024). Even commercial models such as GPT-4 and GPT-4o show signs of the reversal curse (Berglund et al., 2023; Nie et al., 2025b). The reversal curse has been theoretically attributed to an inherent limitation of the autoregressive training objective (Zhu et al., 2024; Kitouni et al., 2024) (see Appendix A.7 for further discussion). Common approaches to mitigating the reversal curse in autoregressive models include: (i) augmenting the training set with paraphrases (Lu et al., 2024), which requires substantial computation to construct; (ii) augmenting the training set with reordered sequences (Guo et al., 2024; Golovneva et al., 2024a), which often violates natural-language grammar and degrades overall language modeling performance; and (iii) replacing causal attention with bidirectional attention (Lv et al., 2024), which struggles to retrieve information over long contexts (e.g., a person’s description). In contrast, our proposed masked fine-tuning paradigm (Section 4.3) for arLLMs addresses the reversal curse without constructing paraphrase augmentations or altering the autoregressive objective; it only requires rewriting the training sample into a demasking task.
2.3 Demasking objective in language models
@@
Recently, dLLMs have emerged as a strong competitor to arLLMs (Sahoo et al., 2024; Nie et al., 2025b; Ye et al., 2025). Compared to autoregressive models, dLLMs use bidirectional attention (non-causal) to generate text by iteratively demasking tokens via a reversed discrete diffusion process. The training objective is to minimize the mask reconstruction loss (Nie et al., 2025b):
@@
$\mathcal{L}(\theta)=-\mathbb{E}_{t,\boldsymbol{x}_{0},\boldsymbol{x}_{t}}\left[\frac{1}{t}\sum_{\ell=1}^{L}\mathbb{I}[\boldsymbol{x}_{t}^{\ell}\in\text{M}]\log p_{\theta}(\boldsymbol{x}_{0}^{\ell}|\boldsymbol{x}_{t})\right],$ (1)
@@
$\boldsymbol{x_{0}}$ is an original sequence of length $L$ sampled from the training data. The masking process is governed by the mask ratio $t$, which is sampled uniformly, resulting in the corrupted sequence $\boldsymbol{x_{t}}$. The set $\mathbf{M}$ denotes the indices of the tokens that were masked by the forward process at ratio $t$. The $\ell$-th token is considered for the loss only if it was masked. Such a loss objective has been shown to be the negative evidence lower bound (ELBO) on the data likelihood (Shi et al., 2024).
@@
Recent studies report dLLMs are more sample-efficient than arLLMs. When the training data is scarce, dLLMs keep improving with repeated use of the data and surpass arLLMs on validation loss, while arLLMs saturate the validation loss or increase it due to overfitting (Prabhudesai et al., 2025; Ni et al., 2025). Prabhudesai et al. (2025) further shows that the lower validation loss in dLLMs can generalize to downstream tasks like ARC-Easy, and attributes its data efficiency to random masks as implicit data augmentation. However, whether these advantages persist in new knowledge acquisition during post-training—where the model needs to learn generalizable knowledge through small fine-tuning sets—remains unclear.
Table 1: Fine-tuning performance of the dLLM and arLLMs across all three datasets. Model names are shortened for presentation (see section 3 for full model names). “Masked” denotes the masked fine-tuning paradigm. “Reverse training” denotes the entity-level reverse-training paradigm proposed in (Golovneva et al., 2024a), which we use as a baseline control. For the Wiki dataset, we use the same-order paraphrase set. Pink indicates clear failure (accuracy below 50%); turquoise indicates clear success (accuracy above 90%).
@@
Masked Llama 8B w/o paraphrases (ours) 0.658 0.949 0.992 0.923 0.971 0.598 0.980 0.930
Masked Llama 8B w paraphrases (ours) 0.969 0.996 0.928 0.832 0.965 0.816 0.905 0.794
@@
Masked Llama 3B w/o paraphrases (ours) 0.887 0.932 0.992 0.933 0.967 0.738 0.970 0.908
Masked Llama 3B w paraphrases (ours) 0.982 0.928 1.000 0.992 0.964 0.809 0.855 0.810
@@
Masked Qwen 7B w/o paraphrases (ours) 0.870 0.897 0.967 0.929 0.984 0.754 0.934 0.896
Masked Qwen 7B w paraphrases (ours) 0.957 0.810 0.933 1.000 0.960 0.828 0.867 0.821
@@
Masked Qwen 4B w/o paraphrases (ours) 0.928 0.902 0.950 0.967 0.944 0.623 0.907 0.870
Masked Qwen 4B w paraphrases (ours) 0.975 0.816 1.000 0.975 0.967 0.806 0.847 0.822
3 Datasets and experimental setups
We focus on assessing LLMs’ ability to learn new knowledge through fine-tuning. More specifically, LLMs are fine-tuned on a set of documents that contain knowledge unknown to the base LLM, and evaluated by open-ended QA tasks. The correctness of an answer is evaluated by the ROUGE-1 score (Lin et al., 2024; Jiang et al., 2025) between the generated answer and the ground truth answer, which is reported as “accuracy” for brevity. It measures the proportion of the words in the ground truth answer that appear in the generated answer. To better demonstrate the generation quality, we also show examples of model responses in all the experiments in A.8.
We use three representative datasets. Two are existing synthetic datasets from prior reversal-curse studies, each accompanied by a paraphrase set for data augmentation; the third is constructed from Wikipedia articles about recent events. Prior work on the reversal curse has rarely used realistic datasets. We include the Wikipedia dataset to probe the reversal curse in a setting that better reflects real-world knowledge acquisition, where new events continually occur and models must absorb and retain them. See examples of each dataset in Appendix A.3.
The NameDescription (Berglund et al., 2023) contains 60 statements of different fictitious individuals, 30 each of the form “[name] is [description]” (N2D) and “[description] is [name]” (D2N). (Lin et al., 2024) extended the dataset with an open-ended QA testing set. For each type of statement, the QA set contains two types of questions: “What is the name related to a given description” and “What is the description of a given name”. Depending on whether the question is aligned with the original statement, each question is classified as “forward” or “backward” question (e.g. N2D statement with “What is the description of a given name” type of question is a forward question). The dataset also contains a paraphrase set, in which each statement is rewritten into 30 different versions, but the order of [name] and [description] in the paraphrases is always preserved as in the original statement (either N2D or D2N).
The Biography dataset is proposed in (Allen-Zhu and Li, 2024, 2025). Since the original dataset is not publicly available, we used a subset of 100 samples from a replication (Zheng et al., 2025). Each sample is a 6-sentence paragraph about a fictitious individual, detailing their birth city, birthday, college, and job information. Note that the name only appears in the first sentence and is replaced with a pronoun in the following sentences; thus, questions about the name are considered backward questions. Each sample also includes a paraphrase set of 5 paraphrases; the paraphrases do not change the order of the sentences but only alter the wording while preserving the information. The testing QA set has both forward (i.e., asking for an attribute given the name) and backward styles (i.e., asking for the name given 3 attributes from the person) questions.
The Wiki dataset contains 94 Wikipedia articles, constructed according to the protocol described in (Pan et al., 2025). We crawl the Wikipedia pages under the category “2025 by month”, and then further filter out the pages that were created before the year 2025. This procedure ensures that these real-world events are recent enough that both dLLM and arLLM models have minimal knowledge about them, which is justified by the models’ accuracy before fine-tuning (Table 1). For each wiki article, we use GPT-o3-mini to generate QA pairs in both forward and backward styles. By prompting GPT-o3-mini, we construct two different paraphrase sets: one that retains the information in place while only changing the wording (same-order paraphrases); the other also changes the order of information in the article (permute-order paraphrases). 10 paraphrases of each type are generated for every wiki article. All LLM generated QAs or paraphrases are cross-checked by human and invalid entries are filtered out and replaced. More details on constructing the datasets are provided in Appendix A.3.
We chose the arLLMs Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Grattafiori et al., 2024), Qwen2.5-7B-Instruct, Qwen3-4B-Instruct-2507 (Yang et al., 2025) and dLLM LLaDA-8B-Instruct (Nie et al., 2025b) to conduct the experiments. LLaDA-8B-Instruct is directly comparable to Llama-3.1-8B-Instruct as they have similar parameter size and comparable general capability benchmarks. Fine-tuning and evaluation configurations are provided in Appendix A.4.
4 Results
4.1 Failures of arLLM knowledge injection